3. The bacterium Corynebacterium glutamicum has been aerobically cultured in a batch fermenter vessel on glucose at pH 7.0 and 20 °C to produce Glutamic acid for the manufacture of Mono Sodium Glutamate, a flavour compound for the food industry. The time course data for a typical batch culture of this bacterium in a fermenter vessel are given below:
t (h) Cell dry weight (g l-1) Glucose concentration (g l-1) Glutamic acid concentration (g l-1)
0            1.22                                              100                                   0
8              1.35                                               97.0                                 0.5
18            2.72                                               92.6                                 5.5
28            7.39                                              75.0                                 15.0
40             25.8                                              32.0                                  34.0
45           42.52                                             20.5                                 40.0
50           73.70                                              10.3                                 45.0
60            90.01                                             2.1                                    48.0
70            99.48                                            0.0                                        48.5
How many production scale fermenters, each of 100 m3 operating liquid volume, would be needed to make 15,000 tonnes per year of Glutamic acid? Each batch vessel is operated for the harvest time of complete carbon substrate utilisation (70 hours), and takes an additional 5 hours to load with liquid medium, to sterilise, and to inoculate with bacterial culture, and 1 hour to drain after the aerobic batch fermentation. Take a process year to be 320 days.
[2 marks]
Answer #1

The calculations are listed in phase by phase manner below:

First phase calculations:

The total amount of Glutamic acid produced in 100 m3 operating volume container

1m3 = 1000 liter

So, 100 m3 container holds 1 x 105 liters of media.

From 1 liter of media 48.5 gms of glutamic acid is produced. So, from 1 x 105 liters of media 4.85 x 104 gms of glutamic acid is produced which is equal to 48.5kg.

Therefore 100m3 operating volume produced 48.5 kgs of glutamic acid for one cycle.

The second phase of calculations:

Number of cycles a 100m3 operating volume container have to undergo in one year i.e. 320 days according to the question.

The container operates for 76hours in total for one cycle (70 hours for complete carbon substrate utilization + 5 hours for sterilisation and inoculation + 1 hour for draining).

Number of hours for 320 days - 320 x 24 = 7680 hours.

Number of cycles the container has to undergo in 320 days = 7680/76 = 101.05 equivalent to 101 cycles.

So, a container undergoes 101 cycles in 320 days.

The amount of glutamic acid produced for 101 cycles by a 100m3 container is = 101 x 48.5 = 4898.5kg equal to 4.89 tonnes.

Therefore, from the calculations it is shown that a 100m3 container produces 4.89 tonnes in a year or 320 days.

Number of containers needed to produced 15000tonnes of glutamic acid in a year = 15000/4.89 = 3067.48 equal to 3068 containers.

In conclusion, 3068 production scale fermenters are needed for producing 15000tonnes of glutamic acid.